# ELSPR: Evaluator LLM Training Data Self-Purification on Non-Transitive Preferences via Tournament Graph Reconstruction

### v1

@@ Appendix E Algorithm Detail
2:Cleaned dataset with transitive preference relations
3:Decompose $G$ into strongly connected components $SCCs$: $\{SCC_{1},SCC_{2},\dots,SCC_{n}\}$
@@
26:for each data point in original dataset do
27:     if preference relation matches $\mathcal{J}$ then
28:         Add data point to Cleaned
29:     end if
30:end for
31:return Cleaned
@@
Table 5: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs( using the CoT Comparison (Tie Allowed) prompt template). Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw, Qwen-Random, and Qwen-Cleaned denote models fine-tuned on the "Raw" and "Cleaned" training sets, respectively. For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base data.
Model Helpful_Base Vicuna Oasst Koala Self-instruct
Auto-J-13B 51.50% 0.7049 63.57% 0.7765 54.86% 0.7101 53.94% 0.7213 58.90% 0.7371
LLaMA-Base 76.41% 0.8518 66.96% 0.8461 67.40% 0.7933 69.87% 0.8180 71.03% 0.8091
LLaMA-Raw 59.03% 0.7654 60.18% 0.7718 58.21% 0.7595 57.22% 0.7457 59.95% 0.7830
LLaMA-Cleaned (ours) 40.20% 0.6523 45.36% 0.6910 43.01% 0.6288 44.38% 0.6591 41.95% 0.6423
Table 6: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. LLaMA-Base refers to the original LLaMA3.1-8B-Instruct model, while LLaMA-Raw, LLaMA-Random, and LLaMA-Cleaned denote models fine-tuned on the "Raw" and "Cleaned" training sets, respectively. For example, LLaMA-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the "Cleaned" training set derived from filtered Helpful_Base data. Auto-J 13B is fine-tuned based on the LLaMA-2-13B-Chat model.
@@
Table 7: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. Qwen-Base denotes to the original Qwen2.5-7B-Instruct model. Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively. Specifically, Qwen-Helpful_Base-Raw is fine-tuned on the "Raw" training data generated from the Helpful_Base dataset, while Qwen-Helpful_Base-Cleaned is fine-tuned on the corresponding "Cleaned" dataset after removing non-transitive preference data.
@@
Table 8: Comparison of Preference Non-Transitivity and Overall Clarity for Evaluator LLMs. LLaMA-Base refers to the original LLaMA3.1-8B-Instruct model. Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively. Specifically, LLaMA-Helpful_Base-Raw is fine-tuned on the "Raw" training data generated from the Helpful_Base dataset, while LLaMA-Helpful_Base-Cleaned is fine-tuned on the corresponding "Cleaned" dataset after removing non-transitive preference data.
@@
"model_identifier": "m",
"output": """{output_1}"""
"model_identifier": "M",
"output": """{output_2}"""
@@
## Your answer: "Concise explanation" followed by "Which is best, m or M?"
@@
"model_identifier": "m",
"output": """{output_1}"""
"model_identifier": "M",
"output": """{output_2}"""
@@
## Your answer: "Concise explanation" followed by "Which is best, m, M or D?"

### v3

@@ Appendix A Algorithm Detail
2:Clean dataset with transitive preference relations
3:Decompose $G$ into strongly connected components SCCsSCCs: $\{SCC_{1},SCC_{2},\dots,SCC_{n}\}$
@@
26:Initialize empty dataset Discarded
27:for each data point in original dataset do
28:  if preference relation matches $\mathcal{J}$ then
29:   Add data point to Cleaned
30:  else
31:   Add data point to Discarded
32:  end if
33:end for
34:return Cleaned, Discarded
Appendix B Additional Experimental Results
B.1 Experimental Results of Different Prompt Forms.
Across five testing sets, models fine-tuned on the “Cleaned” training set exhibit the highest preference clarity. In four of these testing sets (excluding Vicuna), these models also demonstrate the lowest preference non-transitivity, outperforming models fine-tuned on the “Raw” training set and the teacher model Qwen2.5-Max. We observe that models fine-tuned on the “Raw” training set, as well as the teacher model Qwen2.5-Max, exhibit a higher rate of ties in preferences on the Vicuna testing set. The prevalence of such ties—often arising from SCCs formed by cases like $(A=B)$, $(B=C)$, and (C=A)(C=A)—significantly reduces the degree of preference non-transitivity when these SCCs are removed. As shown in Table 8, we compare the probability of ties among three models on the Vicuna testing set. Furthermore, the $\tau_{\text{avg}}^{\downarrow}$ values of models fine-tuned on the “Cleaned” training set indicate a higher overall clarity in preferences.
Figure 5: Comparison of Data Volumes Between “Raw” and “Cleaned” Training Sets Across Different Datasets (Using the CoT Comparison (Tie Allowed) Prompt Template)
Model tie probability
Qwen2.5-Max 7.59%
Qwen-Raw 10.06%
Qwen-Cleaned 6.13%
Table 8: Comparison of tie probabilities in evaluation results produced by different evaluator LLMs on Vicuna testing set.
B.2 “Unseen” question validation
Appendix C Evaluate Prompts
C.1 CoT Comparison
The CoT Comparison prompt template, identical to AlpacaEval 2.0, is detailed in Table 12.
C.2 CoT Comparison (Tie Allowed)
The CoT Comparison (Tie Allowed) prompt template, is detailed in Table 13.
Appendix D Limitations and Future Work
Although we have validated the effectiveness of our filtering approach on five datasets from AlpacaEval, it still cannot fully cover all real-world scenarios, and further validation across a broader range of domains is needed. Additionally, while our method has successfully reduced the issue of preference non-transitivity, it has not been completely eliminated. Further efforts to minimize preference intransitivity remain an important direction for future work. Lastly, our study primarily focused on pairwise comparison tasks in single-turn dialogues, exploring multi-turn dialogue tasks presents an intriguing avenue for future research.
Appendix E Ethics Statement
This research aims to analyze and mitigate non-transitive preferences in large language models when employed as evaluators for open-ended tasks. Our methodology exclusively utilizes publicly available data, existing datasets, and open-source frameworks for computational experiments. All human annotation revisions were conducted by graduate students under fair working conditions with appropriate compensation, ensuring both high-quality data collection and adherence to ethical standards. We confirm that our research procedures comply with established ethical guidelines for AI research and do not present foreseeable negative societal impacts.
@@ Appendix E Ethics Statement
Table 9: Comparison of Preference Non-Transitivity and Overall Clarity for evaluator LLMs ( using the CoT Comparison (Tie Allowed) prompt template). Qwen-Base denotes to the original Qwen2.5-7B-Instruct model, while Qwen-Raw and Qwen-Cleaned denote models fine-tuned on the “Raw” and “Cleaned” training sets, respectively. For example, Qwen-Cleaned in the Helpful_Base column reflects the performance of the model fine-tuned on the “Cleaned” training set derived from filtered Helpful_Base data.
@@
Table 10: Comparison of Preference Non-Transitivity and Overall Clarity for evaluator LLMs. Qwen-Base denotes to the original Qwen2.5-7B-Instruct model. Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively. Specifically, Qwen-Helpful_Base-Raw is fine-tuned on the “Raw” training data generated from the Helpful_Base dataset, while Qwen-Helpful_Base-Cleaned is fine-tuned on the corresponding “Cleaned” dataset after removing non-transitive preference data.
@@
Table 11: Comparison of Preference Non-Transitivity and Overall Clarity for evaluator LLMs. LLaMA-Base refers to the original LLaMA3.1-8B-Instruct model. Raw group and Cleaned group refer to models fine-tuned on the original and filtered training sets, respectively. Specifically, LLaMA-Helpful_Base-Raw is fine-tuned on the “Raw” training data generated from the Helpful_Base dataset, while LLaMA-Helpful_Base-Cleaned is fine-tuned on the corresponding “Cleaned” dataset after removing non-transitive preference data.
@@
”model_identifier”: ”m”,
”output”: ”””{output_1}”””
”model_identifier”: ”M”,
”output”: ”””{output_2}”””
@@
## Your answer: ”Concise explanation” followed by ”Which is best, m or M?”
@@
”model_identifier”: ”m”,
”output”: ”””{output_1}”””
”model_identifier”: ”M”,
”output”: ”””{output_2}”””
@@
## Your answer: ”Concise explanation” followed by ”Which is best, m, M or D?”